Find the perpendicular distance from point (1,-1,4) to the plane 2x-3y-2z=7


1
Expert's answer
2021-08-11T19:28:48-0400
"plane:2x-3y-2z-7=0"

"Point\ A(1, -1, 4)"

The perpendicular distance from "Point\ A" to the given plane is


"d=\dfrac{|2(1)+(-3)(-1)+(-2)(4)-7|}{\sqrt{(2)^2+(-3)^2+(-2)^2}}=\dfrac{10}{\sqrt{17}}"

"d=\dfrac{10}{\sqrt{17}}"